What is the 4th term in this sequence?
1, 4, 27, ?, 3125

Answer by jim_thompson5910(35256)

1 = 1^1

4 = 2^2

27 = 3^3

See the pattern?

The next term would be

256 = 4^4

After that is

3125 = 5^5

So everything fits.